Find (if possible) the supplement of the following angle.
π / 10

Answer by lwsshak3(11628)

Find (if possible) the supplement of the following angle.
π / 10
let x=supplement of given angle
given angle and its supplement add up to π radians
π/10+x=π
x=π-π/10=9π/10 or 0.9π radians